### v1

# Prospects for PBR detection of KM3-230213A-like events

Abstract
POEMMA-Balloon with Radio (PBR) is a scaled-down version of the Probe Of Extreme Multi-Messenger Astrophysics (POEMMA) design, optimized to be flown as a payload on one of NASA’s sub-orbital super pressure balloons circling the Earth above the southern oceans for a mission duration of more than 20 days. One of the main science objectives of PBR is to follow up astrophysical event alerts in search of neutrinos with very high energy ($10^{8}\lesssim E_{\nu}/{\rm GeV}\lesssim 10^{10}$). Of particular interest for anticipated PBR observations, the KM3NeT Collaboration has recently reported the detection of the neutrino KM3-230213A with $10^{8.1}\lesssim E_{\nu}/{\rm GeV}\lesssim 10^{8.9}$. Such an unprecedented event is in tension with upper limits on the cosmic neutrino flux from IceCube and the Pierre Auger Observatory: for a diffuse isotropic neutrino flux there is a $3.5\sigma$ tension between KM3NeT and IceCube measurements, and about $2.6\sigma$ if the neutrino flux originates in transient sources. Therefore, if KM3-230213A was not beginner’s luck, it becomes compelling to consider beyond Standard Model (BSM) possibilities which could lead to a signal at KM3NeT-ARCA but not at IceCube/Auger. We calculate the PBR horizon-range sensitivity to probe BSM physics compatible with observation at KM3NeT-ARCA and non-observation at IceCube/Auger. As an illustration, we consider a particular class of BSM physics models which has been described in the literature as a possible explanation of KM3-230213A.
1 Introduction
On February 13, 2023, a partial implementation of the KM3NeT/ARCA detector [2, 3] recorded an energetic, nearly horizontal muon track, with a reconstructed energy of $E_{\mu}=120^{+110}_{-60}\,\mathrm{PeV}$ [4]. This event, designated as KM3-230213A, most likely was initiated by a neutrino interaction. It represents the highest-energy neutrino event observed to date. To optimally study the possible origin of such an extreme energy neutrino, an instrument is desired that is capable of pointing at any position in the sky, following the source for a sufficient time, and having a low energy threshold.
POEMMA-Balloon with Radio (PBR) is a stratospheric balloon mission [5] representing the most advanced pathfinder for future space-based missions, such as the Probe of Multi-Messenger Astrophysics (POEMMA) [6]. PBR is designed to fulfill these above requirements as one of its three main scientific objectives: the search for transient astrophysical neutrinos via a Target-of-Opportunity (ToO) approach following multi-messenger event alerts. In addition, its scientific program includes the observation of high-altitude horizontal air-showers (HAHAs) and the inaugural measurement of the fluorescence emissions from extensive air showers (EASs) generated by ultra-high-energy cosmic rays (UHECRs) from a suborbital platform.
As a pathfinder, PBR builds on the design studies of POEMMA and leverages experience from previous balloon missions, including EUSO-SPB1 [7] and EUSO-SPB2 [8]. The payload integrates three primary instruments: a Fluorescence Camera (FC), a Cherenkov Camera (CC), and a dedicated Radio Instrument (RI). The FC, which records the longitudinal profiles of UHECR-induced air showers, is described in detail elsewhere [9]. The CC is incorporated into the hybrid focal surface of a 1.1 m Schmidt telescope with a field of view (FoV) of 36 (vertical) by 24 (horizontal). Together with the RI, which complements the optical measurements, these instruments form the principal detectors for neutrino signatures such as those associated with event KM3-230213A. A NASA-supplied rotator paired with a custom-designed tilt system facilitates precise positioning of both instruments to any location below the horizon, a critical feature for following up on multi-messenger alerts.
The Cherenkov Camera is designed to capture ultra-fast optical pulses, operating with an integration time of only 10 ns. It uses silicon photomultipliers (SiPMs) with a pixel size of 3 mm×$\times$3 mm, yielding an instantaneous FoV of $0.2^{\circ}$ per pixel. The 2048-pixel camera, with a detection wavelength range from 320–900 nm provides a total FoV of $6^{\circ}$ (vertical) by $12^{\circ}$ (horizontal). This configuration is sufficiently wide to monitor an astrophysical event for over 20 minutes without active tracking. The estimated energy threshold is approximately 500 TeV.
The Radio Instrument (RI) consists of two sinusoidal radio antennas with a broadband gain between 50–500 MHz, based on the design of the Low Frequency instrument for the Payload for Ultra-High Energy (PUEO) experiment and optimized for EAS detection [10]. Mounted beneath the telescope, these antennas cover the entire optical FoV with 5dB points of ±$\pm 30^{\circ}$ from boresight, and are externally triggered by the Cherenkov Camera.
PBR is scheduled to fly as a payload on a NASA Super Pressure Balloon (SPB), which will circumnavigate the Southern Ocean for more than 20 days following its launch from Wanaka, New Zealand, in Spring 2027.
In this paper we calculate the PBR horizon-range sensitivity to detect KM3-230213A-like events. The layout is as follows. In Sec. 2 we rexamine the main characteristics of the KM3-230213A event, in Sec. 3 we explore potential interpretations of KM3-230213A that can be investigated with PBR, and in Sec. 4 we evaluate the sensitivity of PBR to such phenomena.
2 The KM3-230213A event
The highest energy muon event observed by KM3NeT/ARCA during a 284.7 day observation period (using 21 detection strings) recorded muon energy of $120^{+110}_{-60}$ PeV, with a 90% CL interval of 35–380 PeV [4]. The muon’s trajectory was reconstructed to $0.6^{\circ}$ above the horizon and at $259.8^{\circ}$ in azimuth, with a directional uncertainty is $1.5^{\circ}$ (68% CL). This uncertainty is anticipated to improve to $0.12^{\circ}$ once detector positioning is more precicely determined. The event is labeled KM3-230213A, with event coordinate central values of ${\rm RA}=94.3^{\circ}$ and ${\rm DEC}=-7.8^{\circ}$, and MJD$=59988.0533299$. If the event comes from a steady isotropic neutrino flux, its scaled per flavor flux is $E_{\nu}^{2}\Phi(E_{\nu})=5.8^{+10.1}_{-3.7}\times 10^{-8}$ GeV/(cm2s sr) [4], shown by the red data point in fig. 1.
The KM3NeT collaboration has ruled out atmospheric muons as the origin of the event. Their analysis used MCEq software [11] to model atmospheric particle cascades and calculate sea-level muon fluxes, employing the SIBYLL 2.3c hadronic interaction model. The solid curves in fig. 1 show the sea-level flux of atmospheric muons scaled by $E_{\mu}^{2}$ evaluated using MCEq with the H3p cosmic ray spectrum and composition [12]. For reference, the dashed lines show the atmospheric flux of $\nu_{\mu}+\bar{\nu}_{\mu}$. The lower and upper pairs of curves are for vertical ($\theta=0^{\circ}$) and horizontal ($\theta=90^{\circ}$) directions, respectively. Above muon energies of $\sim 10$ PeV, the atmospheric muon flux is nearly isotropic since the principal contributions come from prompt decays of charm hadrons and light unflavored mesons [13, 14].
[…]
Muons at sea level lose energy in transit to the detector. The KM3NeT collaboration reports a 300 kmwe column depth from sea level to detector for the most likely track trajectory direction of $0.6^{\circ}$ above the horizon. They report that the trajectory column depth from sea level to the detector would be $\sim 60$ kmwe (all water) if its direction is $2.6^{\circ}$ above the horizon. A conservative estimate of muon energy loss from the average energy loss formula
$\Biggl{\langle}\frac{dE_{\mu}}{dX}\Biggr{\rangle}\simeq-bE_{\mu}$ (1)
can be obtained using $b=3.5\times 10^{-6}$cm2/g [15, 16]. For the shorter column depth of $X=60$ kmwe, $E_{\mu}^{f}\simeq E_{\mu}^{i}\exp(-bX)=35$ PeV requires an initial energy of $E_{\mu}^{i}\simeq 4.6\times 10^{16}$ GeV ($4.6\times 10^{10}$ PeV).
Muon electromagnetic energy loss can be understood more precisely by including stochastic effects. Using the ALLM photonuclear energy loss formula [17, 18] along with bremsstrahlung and pair production processes [19], one can find the muon survival probability [20, 21] as a function of distance in water. This is plotted in the left panel of fig. 2 for three initial muon energies and a final muon energy greater than or equal to 35 PeV. The right panel shows the same muon survival probability on a log scale to illustrate the rapid fall-off of the survival probability as a function of distance for initial muon energies in the range of $5\times 10^{8}-5\times 10^{10}$ GeV and $E_{\mu}^{f}=35$ PeV. Even for an initial muon energy of $5\times 10^{10}$ GeV and $E_{\mu}^{f}=35$ PeV, the survival probability is less than $10^{-5}$ for $X\simeq 25$ kmwe, and negligible for 60 kmwe. The probability for a lower energy muon to survive a distance of 60 kmwe with an energy of $E_{\mu}^{f}=35$ PeV is significantly smaller. As noted by KM3NeT Collaboration, even if the candidate muon energy is at the lower end of the 90% confidence limit of the energy measurement, the atmospheric muon flux is not a candidate source of KM3-230213A.
Figure 1: Flux of atmospheric muons scaled by muon energy $E_{\mu}^{2}$ (solid lines) and atmospheric muon neutrinos scaled by neutrino energy (dashed lines) for zenith angles $0^{\circ}$ and $90^{\circ}$ evaluated using MCEq [11] with Sibyll 2.3c and the H3p cosmic ray flux, and the KM3NeT scaled isotropic per flavor neutrino flux (data point with error bars) [4].
Figure 2: Survival probability of muons as a function of distance traveled in water ($\rho=1$ g/cm3), for three initial energies with a final muon energy of $E_{\mu}^{f}=35$ PeV. The right plot shows the same results with log-linear scales.
3 Possible neutrino origins of KM3-230213A
Having ruled out the possibility that KM3-230213A was engendered by an atmospheric muon, the next step is to investigate the possible sources of cosmic neutrinos. Actually, a neutrino origin of KM3-230213A is also challenging because the Pierre Auger Observatory and IceCube, which have been operating with a much larger effective area than KM3NeT and for a considerably longer time, have not observed any neutrinos with $E_{\nu}\gtrsim 10^{7.6}~{}{\rm GeV}$ [22, 23]. The absence of extreme energy neutrinos in the Auger [22] or IceCube [23] data samples excludes any association of KM3-230213A with the diffuse cosmogenic neutrino flux. To be more specific, for a diffuse isotropic neutrino flux there is a $3.5\sigma$ tension between KM3NeT-ARCA and IceCube measurements. The tension between KM3NeT-ARCA and IceCube measurements is about $2.6\sigma$ if the neutrino flux originates in a transient source [24, 25]. The combined observations from KM3NeT-ARCA, IceCube, and Auger are consistent with a source flare of duration $T\sim 1~{}{\rm yr}$, with muon neutrino flux $F\sim 2\times 10^{-7}\ (1~{}{\rm yr}/T)~{}{\rm GeV/(cm^{2}s)}$ [26]. The all-sky rate of similar neutrino flaring sources is constrained to be ${\cal R}\lesssim 0.4/{\rm yr}$. Similar results have been reported in [27].
To elucidate the origin of KM3-230213A it is pivotal to examine consistency with multi-messenger observations; particularly, constraints from $\gamma$-ray telescopes. It is of common knowledge that transparent astrophysical sources of extreme energy neutrinos would also produce a bright $\gamma$-ray signal. To date, no convincing evidence of $\gamma$-rays accompanying KM3-230213A has been observed [28]. As a consequence, constraints can be placed on a combination of the source redshift and the intergalactic magnetic field strength between the source and Earth [29, 30]. If the strength of the intergalactic magnetic field were $B\gtrsim 3\times 10^{-13}~{}{\rm G}$, then the $\gamma$-ray emission from farway sources could be attenuated en route to Earth, and the model of transient sources described in [26] would be consistent with observations. On the other hand, the absence of a $\gamma$-ray signal associated to KM3-230213A challenges scenarios in which this neutrino has a Galactic origin, e.g., from the decay of super-heavy dark matter clustered in the halo. Nevertheless, it was pointed out in [31] that if the dark matter were just ordinary matter locked inside primordial black holes (PBHs), then the KM3NeT-ARCA event might have originated from a Galactic transient explosion of a PBH near the end of its evaporation lifetime. Actually, the $\gamma$-ray constraints could be easily accommodated in scenarios with large extra dimensions, because the emission of higher-dimensional PBHs is dominated by Kaluza-Klein modes [32, 33]. It was recently pointed out in [34] that a cosmological population of microscopic black holes, which rapidly evaporate via Hawking radiation, emitting intense, short-lived bursts of neutrinos with $E_{\nu}\gtrsim 10^{8}{\rm GeV}$ can accommodate the groundbreaking KM3NeT-ARCA detection while remaining consistent with IceCube’s non-detections. These microscopic black holes are expected to be produced via bubble collisions generated in the merger of astrophysical black holes; therefore, exploring another window of multi-messenger astrophysics driven by neutrinos and gravitational waves.
A plethora of models advocating beyond Standard Model (BSM) physics have been proposed to explain the origin of KM3-230213A (we refrain from referring to this literature in detail in these proceedings). Of particular interest for PBR observations are BSM possibilities which could lead to a signal at KM3NeT-ARCA but not at IceCube/Auger. For example, the BSM model conjectured in [35] relies on the fact that the KM3-230213A has traversed approximately 147 km of rock and sea en route to the detector, whereas neutrinos arriving from the same location in the sky would have only traveled through about 14 km of ice before reaching IceCube. This is because the arrival direction of KM3-230213A has an incidence angle of $0.6^{\circ}$ at KM3NeT-ARCA and about $8^{\circ}$ above the horizon at IceCube. The idea proposed in [35] is that a new physics matter potential can resonantly amplify sterile-to-active transitions and accommodate oscillations of extreme energy sterile neutrinos in the 100 km range. The difference in propagation distance can then be used to increase the active neutrino flux near the KM3NeT detector, alleviating the tension between KM3NeT-ARCA and IceCube observations. This BSM model then requires that most of the emission by a point source in the nominal direction of the detected KM3-230213A neutrino arrives at Earth as sterile neutrinos. Two other similar scenarios which make use of the difference in propagation distance are described in [36, 37]. PBR will sample a range of propagation distances. As a point source moves through the FoV of the PBR Cherenkov Camera, particle trajectories through the Earth range up to chord lengths of 2,300 km and column depths up to 7,400 kmwe (see, e.g., fig. 2 in [38]).
In the next section we calculate the horizon-range sensitivity of PBR and show that this stratospheric balloon mission will be able to probe BSM physics compatible with observation at KM3NeT-ARCA and non-observation at IceCube/Auger.
4 PBR discovery reach
The integrated exposure of PBR is limited by observing conditions (e.g., dark skies, presence of clouds), the length of time a given source can be observed, and the overall duration of the balloon flight. Nevertheless, the expected instantaneous acceptance for very-high-energy neutrinos is expected to be comparable with other ground based detectors. Thus, searches of very-high-energy neutrinos with PBR will be especially powerful for transient sources.
To assess PBR’s performance in observing KM3-230213A-like events, we estimate its sensitivity to neutrino bursts. To that end, we calculate the time-averaged $\nu_{\tau}$ acceptance:
[…]
where $\phi$ and $\delta$ are the co-latitude and longitude of the celestial position of a given source (e.g., $\phi$ is the right ascension in the equatorial celestial coordinate system and $\delta$ is the declination), $T_{0}$ is the observation time, and $f(t)$ is an observation efficiency as a function of time that accounts for the reduction in the observation time due to light from the Sun and Moon [39]. In eq. (2), $A(\beta_{\rm tr}(t,\lambda_{\oplus},\phi_{\oplus}),E_{\nu},\phi,\delta)$ is the instantaneous acceptance to $\nu_{\tau}$s from the source as defined by the geometry of the region of interest around the source on the surface of the Earth and the probability of detecting EASs induced by $\tau$ leptons that from the Earth at a selected location within the region (at e.g., Earth latitude, $\lambda_{\oplus}$, and longitude, ϕsubscriptitalic-$\phi_{\oplus}$) and that propagate along quasi-parallel trajectories that point back to the source. The trajectory of the $\tau$ lepton makes an angle $\beta_{\rm tr}$ above the Earth’s surface. We perform Monte Carlo simulations to sample the region of interest and to calculate the detection probabilities of sampled events based on $\nu_{\tau}$ interaction physics in the Earth, $\tau$-lepton decays, EAS development, and the propagation of Cherenkov light signals through the atmosphere to the detector. In calculating time-averaged acceptances, we account for the changes in the source’s position with respect to the Earth and the detector. We also account for model the balloon’s trajectory based on historical wind patterns and assuming a launch on April 6, 2027.
For the calculations we present here, we consider two scenarios: a short, 1000-s burst of a transient source in an optimal observation time window ($f(t)=1$), and a period of neutrino emission lasting longer than $\sim 20$ days (as motivated by typical balloon mission durations). In the short burst scenario, we assume the telescope is optimally pointed to catch the source just as it dips below the horizon at the burst’s initial time $t_{0}$.
To determine the PBR discovery reach for neutrino sources, we adopt the quasi-differential neutrino flux over a decade of energy [40], where the sensitivity can be approximated by the relation
${\rm Sensitivity}=\frac{2.44}{\ln 10}\times\frac{N_{\nu}E_{\nu}}{\langle A(E_{\nu})\rangle_{T_{0}}}\ ,$ (3)
where the factor $N_{\nu}=3$ converts the $\nu_{\tau}$ sensitivity to the all-flavor sensitivity, and we have taken the $90\%$ unified confidence level, assuming no signal and no background [41], viz. $2.44/\ln 10$.
Figure 3: All-flavor 90% CL sensitivities for PBR for optimal observing conditions during a 1000-s ToO observation (left) and over a 20-day period limited to dark-sky observations (right). For comparison, the modeled 1000-s and 14-day fluences for a BSM burst are indicated by orange lines and stars. Also plotted are the upper limits for GW170817 reported by IceCube (red), Auger (blue), and ANTARES (magenta) computed for 1000 s and 14 days (see Ref. [42]).
The energy-squared weighted flux of the transient neutrino source discussed in Sec. 3 is
$E_{\nu}^{2}\ \frac{dN}{dE_{\nu}}\sim 10^{-6}~{}{\rm GeV\ cm^{-2}\ s^{-1}}\,,$ (4)
see Fig. 1 in [26]. In rescaling $dN/dE_{\nu}$ by a factor of $100$ to approximately account for the combination of exposure and detector size of IceCube relative to KM3NeT [32, 27], we obtain
$\left.E_{\nu}^{2}\ \frac{dN}{dE_{\nu}}\right|_{\rm BSM}\sim 10^{-4}~{}{\rm GeV\ cm^{-2}\ s^{-1}}\,.$ (5)
For this flux, we expect a BSM-burst fluence over an observation time of $1000~{}{\rm s}$ to be ${\cal F}_{\rm BSM}\sim 0.1~{}{\rm GeV}\ {\rm cm}^{-2}$ in the energy range of $10^{7.5}\lesssim E_{\nu}/{\rm GeV}\lesssim 10^{9.0}$. For an observation time of 20 days, we expect the fluence to be ${\cal F}_{\rm BSM}\sim 170~{}{\rm GeV}\ {\rm cm}^{-2}$ over this energy range, which is within the projected 90% unified confidence level sensitivity of PBR shown in fig. 3. The population of transient sources proposed in [26] has a cosmological distance scale of 20 Gpc. A precise description of the dependence of the PBR performance in detecting sources at various redshifts will be presented in an upcoming journal publication.
Acknowledgments
The authors would like to acknowledge the support by NASA awards RTOP 21-APRA21-0071, 80NSSC22K1488, and 80NSSC24K1780, by the French space agency CNES and the Italian Space agency ASI. The work is supported by OP JAC financed by ESIF and the MEYS CZ.02.01.01/00/22_008/0004596. The work of L.A.A. and K.P.C. is supported by the US National Science Foundation (grant PHY-2412679). We gratefully acknowledge the collaboration and expert advice provided by the PUEO collaboration. We also acknowledge the invaluable contributions of the administrative and technical staffs at our home institutions.